# Towards Fair RAG: On the Impact of Fair Ranking in Retrieval-Augmented Generation

### v3

@@ E.1. LaMP data statistics for Flan-T5-Small
Table 11. LaMP data statistics for Flan-T5-Small after filtering for fairness evaluation.
E.2. LaMP data statistics for Flan-T5-Base
Dataset #queries Avg # Docs (Std) Avg # Pos Labels (Std) Avg % Pos Labels
@@ E.2. LaMP data statistics for Flan-T5-Base
Table 12. LaMP data statistics for Flan-T5-Base after filtering for fairness evaluation.
E.3. LaMP data statistics for Flan-T5-XXL
Dataset #queries Avg # Docs (Std) Avg # Pos Labels (Std) Avg % Pos Labels
@@ E.3. LaMP data statistics for Flan-T5-XXL
Table 13. LaMP data statistics for Flan-T5-XXL after filtering for fairness evaluation.
E.4. LaMP data statistics for Flan-UL2
Dataset #queries Avg # Docs (Std) Avg # Pos Labels (Std) Avg % Pos Labels
@@ E.4. LaMP data statistics for Flan-UL2
Table 14. LaMP data statistics for Flan-UL2 after filtering for fairness evaluation.

### v4

@@ D.1. LaMP corpus statistics for Flan-T5-Small
Table 4. LaMP corpus statistics for Flan-T5-Small after filtering for fairness evaluation.
D.2. LaMP corpus statistics for Flan-T5-Base
Corpus #queries Avg # Docs (Std) Avg # Useful Docs (Std) Avg % Useful Docs
@@ D.2. LaMP corpus statistics for Flan-T5-Base
Table 5. LaMP corpus statistics for Flan-T5-Base after filtering for fairness evaluation.
D.3. LaMP corpus statistics for Flan-T5-XXL
Corpus #queries Avg # Docs (Std) Avg # Useful Docs (Std) Avg % Useful Docs
@@ D.3. LaMP corpus statistics for Flan-T5-XXL
Table 6. LaMP corpus statistics for Flan-T5-XXL after filtering for fairness evaluation.
D.4. LaMP corpus statistics for Flan-UL2
Corpus #queries Avg # Docs (Std) Avg # Useful Docs (Std) Avg % Useful Docs
@@ D.4. LaMP corpus statistics for Flan-UL2
Table 7. LaMP corpus statistics for Flan-UL2 after filtering for fairness evaluation.